# SynLlama: Generating Synthesizable Molecules and Their Analogs with Large Language Models

### v1

@@ Abstract
Generative machine learning models for small molecule drug discovery have shown immense promise, but many molecules generated by this approach are too difficult to synthesize to be worth further investigation and/or development. We present a novel approach by fine-tuning Meta’s Llama3 large language models (LLMs) to create SynLlama, which generates full synthetic pathways made of commonly accessible building blocks and robust organic reaction templates. SynLlama explores a large synthesizable space using significantly less data compared to other state-of-the-art methods, and offers strong performance in bottom-up synthesis, synthesizable analog generation, and hit expansion, offering medicinal chemists a valuable tool for drug discovery developments. We find that SynLlama, even without training on external building blocks, can effectively generalize to unseen yet purchasable building blocks, meaning that its reconstruction capabilities extend to a broader synthesizable chemical space than the training data.
@@
4Department of Chemistry, University of Minnesota, 207 Pleasant Street SE, Minneapolis, MN 55455, USA 5Contramont Research, San Francisco, CA, 94158 USA
@@ 1 Introduction
Drug discovery campaigns have led to the development of pharmaceuticals that address a variety of disease conditions, and which have been shown to be a significant factor in increasing human life expectancy over many decades. 1, 2, 3, 4, 5, 6 In order to better optimize the drug discovery pipeline for new small molecule therapeutics, early screening steps require the construction of large molecular libraries, such as DNA-encoded libraries7 or the Enamine libraries8, that allow for high-throughput experimental and virtual screening of drug binders to desirable targets. To illustrate, the Enamine REAL library is comprised of 38 billion drug-like compounds that take advantage of a large set of optimized reactions and building blocks that allow for rapid and reliable synthesis with over 85% success rate.8 DNA-encoded libraries of drug-like molecules9, with upper bounds of 100 billion molecules, define a unique DNA sequence that acts as a “bar code” defining its synthesis pathway. Given these large libraries high throughput screening methods are then enabled to isolate the most tightly binding molecules to valued targets using experimental or virtual screening.10, 11, 12, 13, 14 However these screened compound collections are often unsuccessful in generating active compounds to some drug targets because they lack sufficient chemical diversity in the screening libraries. In fact an enumeration of possible drug candidates has been estimated to be enormous - a chemical space of $\sim$1060 drug-like molecular compounds.15
@@
With recent advances in artificial intelligence and deep learning, generative models have started filling in the void for expanding to greater drug-like molecular diversity. After training on databases containing various ligand representations16, 17, 18, 19, 20, machine learning models, including string-based 1D generative models and structure-aware 3D de novo methods, have paved the way for quick exploration of greater swaths of unseen chemical space21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38. However, even with their exceptional generative capabilities, these models still face one major challenge: their proposed de novo molecules lack practical guarantees of synthesizability, which limits their utility to medicinal chemists and drug discovery in general 39, 40, 41. For generative approaches in small-molecule drug discovery to fulfill their vast potential, ensuring synthetic feasibility is essential to bridge the gap between in silico drug design and the realistic applicability of computationally generated molecules.
@@
Efforts have been made to address the problem of poor synthesizability of de-novo-generated molecules. One line of research focuses on integrating empirical scoring functions, such as the SA score 42, into the objective functions of learning algorithms. However most SA scores are unreliable and can lead to generation of unsynthesizable molecules because the scoring function only relies on identifying common fragments in molecules43. These scoring functions also assign bad scores to complex yet synthesizable molecules that require multi-step synthetic pathways, which causes generative models to miss viable candidates 44.
@@
Alternatively, proposing molecular candidates using commercially available building blocks and commonly known organic reaction templates through forward synthesis offers enhanced synthetic tractability. Importantly, this strategy is appealing to medicinal chemists, since it offers realistic and actionable synthesis pathways for them to examine and execute. Some models in this direction apply rule-based synthesis and optimization on building blocks alone to generate novel molecules with desired chemical properties45, 46, 47. Other models condition on input molecules to propose synthetic pathways using commercially available building blocks and well-validated reaction templates. The resulting molecules from proposed syntheses can either lead to full construction of the target molecule or the generation of structurally similar analogs within the predefined chemical search space. For example, SynNet48 constructs synthetic trees via Markov Decision Processes MDPs that use multilayer perceptrons to choose the next action space, but it struggles at planning synthesis for previously unseen molecules. More recent models such as ChemProjector49 and Synformer50 use transformers to decode for the next action space and have achieved good empirical performances. But the primary drawbacks of these ML models is that the training requires several hundred million synthetic pathway data, and is limited to the chemical space defined by the building blocks and reaction templates that limits possible synthesis pathways to novel molecules.
@@
While current methods of generating synthesizable molecules through sequences of reactions with commercially available building blocks show promise, their requirements of a significant amount of reaction data and a defined chemical space have directed our attention towards large language models (LLMs). These models offer a compelling alternative due to their foundational nature and adaptability to downstream tasks. 51 LLMs inherently possess extensive chemical knowledge, and recent advancements have focused on extracting and applying this knowledge for predictive and optimization tasks using natural language guidance52, 53, 54, 55. Furthermore, after some supervised fine-tuning, LLM models can perform at the same level as chemical language models trained solely on chemical representations, all while requiring less data.38 The efficiency gained from fine-tuning LLMs thus motivates us to explore their potential in more complex tasks, such as forward synthesis planning, which could pave the way for more cost-effective developments in drug discovery.
@@
Herein, we present SynLlama, an LLM-based tool using the open-source Llama-3.1-8B and Llama-3.2-1B foundation models  56 to deduce synthetic routes for target molecules or closely related analog molecules. SynLlama proposes the necessary reaction (RXN) sequences and commercially available building blocks (BBs) required for performing the synthesis, presenting molecules within a diverse yet synthesizable chemical space. Though trained on nearly one to two orders of magnitude less data than the current state-of-the-art models, SynLlama demonstrates competitive performances as other models in relevant tasks for drug discovery, including synthesis planning for target and analog molecules of pharmaceutical interest. Moreover, because of its generative nature, SynLlama has the added ability to explore commercially available building blocks beyond the predefined synthetic space introduced at the training stage, which previous models are unable to explore.
2 Overview of SynLlama Workflow
@@
The SynLlama workflow illustrated in Figure 1 facilitates the creation of more synthesizable compounds in an expanded chemical space, offering a strategy to reduce synthesis complexity while preserving the desired molecule or finding a synthesizable analog with good molecular similarity to the target. Although general-purpose LLMs such as the Llama 3 models 56, are known for their ability to perform well on a diverse set of chemistry tasks57, 58, there are three essential components to convert them into expert models that can generate synthetic pathway for a related molecule given a target molecule: 1) a reliable and diverse set of reaction data that covers a large synthesizable chemical space, 2) an efficient supervised fine-tuning (SFT) strategy to train a general-purpose LLM on this reaction data, and 3) a reconstruction algorithm that can transform the output of the fine-tuned LLM into valid synthesis routes, ensuring the final products lie within a commercially available chemical search space. In the following sections, we detail our methods and procedures for achieving the three goals to yield the SynLlama models.
@@ 2.1 Reaction Data Generation
Figure 1: Overview of the SynLlama workflow including data generation, supervised fine-tuning, inference, and reconstruction. (a) The predefined synthesizable chemical space of reaction templates (RXN) and building blocks (BBs) that covers billions of molecules. (b) An example synthesis data and its generation process from the defined synthesizable chemical space to create training examples. Here, RXN 76 represents amide coupling and RXN 72 represents Suzuki coupling. (c) A schematic representation of supervised fine-tuning that converts Llama 3 models to SynLlama models, along with the instruction, input, and output for the example synthesis in (b). (d) SynLlama’s inference on an unseen test molecule. Black represents SynLlama’s raw retrosynthetic output consisting of RXN sequences and predicted BBs, while colored BBs indicate the top two most similar BBs to the predicted ones from the Enamine building block library. Here, RXN 76 represents amide coupling and RXN 72 represents Suzuki coupling. (e) Reconstructed molecules using the predicted reaction sequences and similar building blocks from the Enamine building block library. In this example, all predicted building blocks are present in the Enamine library, allowing for the complete reconstruction of the input molecule and the generation of close analogs.
@@ 2.2 Supervised Fine-tuning and Inference from SynLlama
To create the SynLlama model, we need to establish data generation protocols for supervised fine-tuning (SFT) of the Llama 3 models as shown schematically in Figure 1(c). When generating reaction data in text format, we choose to represent the BBs and intermediates along the synthetic pathway using SMILES61 strings, while RXNs are explicitly defined in the SMARTS62 format. These structured chemical notations aim to enhance SynLlama’s ability to systematically identify and deconstruct bonds according to RXN templates, effectively dismantling input molecules into building-block-sized fragments. For defining the training BBs, we calculate the Morgan fingerprints 63 of all BBs and employ KNN clustering to create 128 groups. We then randomly withhold one group as the test set, resulting in $\sim$228,000 BBs for training and $\sim$1,000 for testing. All reaction templates are accessible to both the training and testing BBs, thereby defining the training and testing chemical spaces. In addition, since our goal is for SynLlama to learn to link molecules with their synthesis routes, our prompt-response pairs are structured according to retrosynthesis, as depicted in Figure 1(c) and shown in detail in Supplementary Figure S1. Such engineered prompts allow the SynLlama model to learn to construct synthesis pathways for the input molecules by inferring sequences of BBs and RXNs, as well as the intermediate steps. While theoretically the model could predict BBs and RXNs without intermediates, we still include them in individual reaction steps in the hope of activating the inherent chemical knowledge in LLMs and enhancing their understanding of synthesis patterns.
@@
After training the SynLlama models, we apply the consistent prompt setup to perform inferences on molecules. For any given molecule, the SynLlama models predict reaction sequences in SMARTS format and generate SMILES strings for all the reactants, products, and BBs for the reactions they predict. During inference time, the instruction to SynLlama remains the same and SMILES strings in the input section are substituted with ones specified by the user. As depicted in Figure S1, the responses of the SynLlama models follow the output structures enforced by the prepared training prompt-response pairs. To be more specific, the output response section consists of two parts: reactions and building blocks. In the ‘reactions’ component, the model sequentially deconstructs the target molecule by breaking bonds in a retrosynthetic manner. At each reaction step, it predicts a reaction template to use, along with the reactants and product of the reaction, continuing until no reactants can have further reactions. Then, in the ‘building blocks’ section, the model compiles all building blocks, reactants from each reaction that are not products from other reactions, from the ‘reaction’ section. A visual representation of the inference process is illustrated in Figure 1(d) in black ink.
@@ 2.3 Reconstruction from Predicted Retrosynthesis
Using the predicted sequence of reactions and BBs from SynLlama responses, we can synthesize the proposed target molecule or close analogs by applying RXNs to the predicted BBs in the inferred order as shown in Figure 1(d) in black ink. In some cases the predicted BBs correspond to known Enamine BBs, thus ensuring that the generated molecules remain within a well-defined chemical space for synthesis. Often however the predicted BBs are previously unseen due to the generative nature of SynLlama. In fact a substantial fraction of SynLlama’s raw outputs represent valid synthesis pathways using BBs outside the Enamine library, but which are purchasable from other suppliers when searched in Molport. Thus, besides generating molecules within the predefined chemical space with purely Enamine BBs, SynLlama’s raw outputs are an alternative strategy for molecule reconstruction. We return to this point in Results.
An alternative strategy is to map predicted BBs from SynLlama to known Enamine BBs that define molecules in a well-defined chemical synthesis space. In this case we employ a nearest neighbor search algorithm to sample Enamine neighboring BBs from the predicted BBs as shown in Figure 1(d) in colored inks. Since the RXN sequences are predicted concurrently with the BBs, our effective search space for the predicted BBs is only BBs that can react through that specific RXN template. This constrained search space not only allows us to ensure the success rate of such forward syntheses but also enables us to effectively explore fractions of the input molecule. We employ both search trees on SMILES string and Morgan fingerprints when dealing with valid molecules and constrain to only SMILES-based tree search when the generated SMILES strings are invalid. Further details are presented in the Supplementary Information.
When constructing the full synthetic pathways, for reactions with multiple possible products, we select the product that most closely matches the predicted product SMILES strings based on string similarity (Figure 1(e)). The algorithm then iteratively builds the synthesis routes, using all the predicted BBs and RXN sequences to generate variations of the original molecule (or fully reconstruct it when possible) within our defined chemical space. Overall, this reconstruction algorithm enables the SynLlama model to function as a generator for synthesizable molecules along with their corresponding synthetic pathways.

### v2

@@ Abstract
Generative machine learning models for small molecule drug discovery have shown immense promise, but many molecules they generate are too difficult to synthesize, making them impractical for further investigation or development. In this work, we present a novel approach by fine-tuning Meta’s Llama3 Large Language Models (LLMs) to create SynLlama, which generates full synthetic pathways made of commonly accessible building blocks and robust organic reaction templates. SynLlama explores a large synthesizable space using significantly less data compared to other state-of-the-art methods, and offers strong performance in bottom-up synthesis, synthesizable analog generation, and hit expansion, offering medicinal chemists a valuable tool for drug discovery developments. We find that SynLlama, even without training on external building blocks, can effectively generalize to unseen yet purchasable building blocks, meaning that its reconstruction capabilities extend to a broader synthesizable chemical space than the training data. We also demonstrate the use of SynLlama in a pharmaceutical context for synthesis planning of analog molecules and hit expansion leads for proposed inhibitors of target proteins.
@@
4Department of Chemistry, University of Minnesota, 207 Pleasant Street SE, Minneapolis, MN 55455, USA
@@ 1 Introduction
Drug discovery campaigns have led to the development of pharmaceuticals that treat a variety of diseases and have significantly contributed to the increase of human life expectancy over the past several decades. 1, 2, 3, 4, 5, 6 In order to better optimize the drug discovery pipeline for new small molecule therapeutics, early screening steps require the construction of large molecular libraries, such as custom designed DNA-encoded libraries7 or the commercially available Enamine libraries8, that allow for high-throughput experimental and virtual screening of drug binders to desirable targets. To illustrate, the Enamine REAL library comprises 38 billion drug-like compounds that take advantage of a large set of optimized reactions and building blocks that allow for rapid and reliable synthesis with over 85% success rate.8 DNA-encoded libraries of drug-like molecules9, with upper bounds of 100 billion molecules, define a unique DNA sequence that acts as a “bar code” defining its synthesis pathway. Therefore, these extensive libraries enable both experimental and virtual screening methods to identify tightly binding molecules for desired targets.10, 11, 12, 13, 14 However, despite their size, these libraries often fail to produce active compounds for some drug targets due to insufficient chemical diversity. In fact, the estimated chemical space of drug-like molecules consists of $\sim$1060 compounds, suggesting that even the largest existing screening collections represent only a minuscule fraction of all possible candidates.15
@@
With recent advances in artificial intelligence and deep learning, generative models have begun to contribute to these efforts. After training on databases containing various ligand representations16, 17, 18, 19, 20, machine learning models, including string-based 1D generative models and structure-aware 3D de novo methods, have paved the way for quick exploration of greater swaths of unseen chemical space21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38. However, even with their exceptional generative capabilities, these models still face one major challenge: their proposed de novo molecules lack practical guarantees of synthesizability, which limits their utility to medicinal chemists and drug discovery in general 39, 40, 41. For generative approaches in small-molecule drug discovery to fulfill their potential, ensuring synthetic feasibility is essential to bridge the gap between in silico drug design and the realistic applicability of computationally generated molecules.
@@
Efforts have been made to address the problem of poor synthesizability of de-novo-generated molecules. One line of research focuses on integrating empirical scoring functions, such as the SA score 42, into the objective functions of learning algorithms. However, optimizing only the SA score can still lead to generation of unsynthesizable molecules because the scoring function only relies on identifying common fragments in molecules43. In addition, these scoring functions also assign bad scores to complex yet synthesizable molecules that require multi-step synthetic pathways, which causes generative models to miss viable candidates 44.
@@
Alternatively, the proposal of molecular candidates using commercially available building blocks and commonly known organic reaction templates through forward synthesis offers synthetic tractability. Importantly, this strategy is appealing to medicinal chemists, since it offers realistic and actionable synthesis pathways for them to examine and execute. Some models in this direction apply rule-based synthesis and optimization on building blocks or entire synthetic pathways to generate novel molecules with desired chemical properties45, 46, 47. Other models condition on input molecules to propose synthetic pathways using commercially available building blocks and well-validated reaction templates. The resulting molecules from proposed syntheses can lead to either the full construction of the target molecule or the generation of structurally similar analogs within the predefined chemical search space. For example, SynNet48 constructs synthetic trees via Markov Decision Processes (MDPs) and uses multilayer perceptrons to choose the next action space, but it struggles in planning synthesis for previously unseen molecules. More recent models such as ChemProjector49 and Synformer50 use transformers to decode for the next action space and have achieved good empirical performances. However, a major drawback of these ML models is their reliance on vast amounts of reaction data (on the order of a hundred million reaction pathways) and a predefined chemical search space dictated by available building blocks and reaction templates. This constraint not only makes training computationally expensive but also limits their ability to propose truly novel synthesis pathways.
@@
Given these challenges, Large Language Models (LLMs) have emerged as a compelling alternative due to their foundational nature and adaptability to downstream tasks. 51 LLMs inherently possess extensive chemical knowledge, and recent advancements have focused on extracting and applying this knowledge for predictive and optimization tasks using natural language guidance52, 53, 54, 55. Furthermore, after some supervised fine-tuning, LLM models can perform at the same level as chemical language models trained solely on chemical representations, all while requiring less data.38 The efficiency gained from fine-tuning LLMs thus motivates us to explore their potential in more complex tasks, such as forward synthesis planning, which could pave the way for more cost-effective developments in drug discovery.
@@
Herein, we present SynLlama, an LLM-based tool built on the open-source Llama-3.1-8B and Llama-3.2-1B foundation models  56 to deduce synthetic routes for target molecules or structurally related analogs. SynLlama predicts the necessary reaction (RXN) sequences and commercially available building blocks (BBs) required for synthesis, presenting molecules within a diverse yet synthesizable chemical space. Despite being trained on nearly one to two orders of magnitude less data than the current state-of-the-art models, SynLlama demonstrates competitive performance in key tasks for drug discovery, including synthesis planning for target and analog molecules of pharmaceutical interest and hit expansion. Moreover, because of its generative nature, SynLlama has the added ability to explore commercially available building blocks beyond the predefined synthetic space introduced during training - an ability that previous models lack. By integrating molecular design with synthetic feasibility, SynLlama represents a step forward in bridging computational chemistry with synthetic chemistry, providing medicinal chemists with actionable and experimentally accessible molecular candidates.
2 Methods
@@
The SynLlama workflow, illustrated in Figure 1, is designed to generate synthesizable compounds within an expanded chemical space. When an input molecule passes through this workflow, it can either be fully reconstructed through valid synthetic pathways, or the workflow will produce a structurally similar yet synthesizable analog along with its synthesis route. To transform general-purpose LLMs, like the Llama 3 models 56, into expert models for synthetic pathways, we use three key components: 1) a reliable and diverse set of reaction data that covers a large synthesizable chemical space, 2) an efficient supervised fine-tuning (SFT) strategy to train a general-purpose LLM on these reaction data, and 3) a reconstruction algorithm that can convert the output of the fine-tuned LLM into valid synthesis routes, ensuring the proposed molecules lie within a commercially available chemical search space. These components are crucial for leveraging LLMs, which are known to perform well in diverse chemistry tasks 57, 58, to specialize in synthetic modeling. The following sections will detail the methods and procedures used to develop the SynLlama models, providing a holistic understanding of the entire process.
@@ 2.1 Reaction Data Generation
Figure 1: Overview of the SynLlama workflow including data generation, supervised fine-tuning, inference, and reconstruction. (a). The predefined synthesizable chemical space of reaction templates (RXN) and building blocks (BBs) that covers billions of molecules. (b). An example synthesis data and its generation process from the defined synthesizable chemical space to create training examples. Here, RXN 76 represents amide coupling and RXN 72 represents Suzuki coupling. (c). A schematic representation of supervised fine-tuning that converts Llama 3 models to SynLlama models, along with the instruction, input, and output for the example synthesis in (b). (d). SynLlama’s inference on an unseen test molecule. Black represents SynLlama’s raw retrosynthetic output consisting of RXN sequences and predicted BBs, while colored BBs indicate the top two most similar BBs to the predicted ones from the Enamine building block library. Here, RXN 76 represents amide coupling and RXN 72 represents Suzuki coupling. (e). Reconstructed molecules using the predicted reaction sequences and similar building blocks from the Enamine building block library. In this example, all predicted building blocks are present in the Enamine library, allowing for the complete reconstruction of the input molecule and the generation of close analogs.
@@ 2.2 Supervised Fine-tuning and Inference from SynLlama
To create the SynLlama model, we need to establish data generation protocols for supervised fine-tuning (SFT) of the Llama 3 models as schematically shown in Figure 1(c). When generating reaction data in text format, we choose to represent the BBs and intermediates along the synthetic pathway using SMILES61 strings, while RXNs are explicitly defined in the SMARTS62 format. These structured chemical notations are designed to enhance SynLlama’s ability to systematically identify and deconstruct bonds according to RXN templates, effectively dismantling input molecules into building-block-sized fragments.
To define our data splits, we apply a time split of the Enamine BBs where we take all BBs from their August 2024 release as the training BBs and all new BBs from their Februray 2025 release that were not in training as the testing BBs. This procedure results in $\sim$230,000 BBs for training and $\sim$13,000 for testing, while all reaction templates are accessible to both sets of BBs, thus defining the training and testing chemical spaces. In addition, since our goal is for SynLlama to learn to link molecules with their synthesis routes, our prompt-response pairs are structured according to retrosynthesis, as depicted in Figure 1(c) and shown in detail in Supplementary Figure S1. Such engineered prompts and responses allow the SynLlama model to learn to construct synthesis pathways for the input molecules by inferring sequences of BBs and RXNs, as well as the intermediate steps. While theoretically the model could predict BBs and RXNs without intermediates, we still include them in individual reaction steps in the hope of activating the inherent chemical knowledge in LLMs and enhancing their understanding of synthesis patterns.
@@
After training the SynLlama models, we apply the consistent prompt setup to perform inferences on molecules. For any given molecule, the SynLlama models predict reaction sequences in SMARTS format and generate SMILES strings for all the reactants, products, and BBs for the reactions they predict. During inference time, the instruction to SynLlama remains the same, and SMILES strings in the input section are substituted with ones specified by the user. As depicted in Figure S1, the responses of the SynLlama models follow the output structures enforced by the prepared training prompt-response pairs. To be more specific, the output response section consists of two parts: reactions and building blocks. In the ‘reactions’ component, the model sequentially deconstructs the target molecule by breaking bonds using provided reaction templates in a retrosynthetic manner. At each step, it predicts a reaction template, along with the reactants and product of the reaction, continuing until no further reactions are possible. Then, in the ‘building blocks’ section, the model compiles all building blocks, namely, reactants from each reaction that are not products in other reactions, identified from the ‘reaction’ section. A visual representation of the inference process is illustrated in black ink in Figure 1(d).
@@ 2.3 Reconstruction from Predicted Retrosynthesis
Using the predicted sequence of RXNs and BBs from SynLlama responses, we can synthesize the proposed target molecule or close analogs by applying the predicted reaction templates to the BBs in the inferred order, as shown in black ink in Figure 1(d). In some cases, the predicted BBs match known Enamine BBs, ensuring that the resulting molecules remain within an established chemical space for synthesis. However, due to SynLlama’s generative nature, some predicted BBs are novel. As detailed in Table S2, some of SynLlama’s outputs provide valid synthesis pathways involving at least one BB not found in the Enamine library, but these BBs can be purchased from other suppliers identified by Molport 63. Therefore, while SynLlama primarily produces molecules within the predefined chemical space using Enamine BBs, its output also offers an alternative strategy for molecule construction. We will revisit this point in the Results section.
When the input molecule cannot be fully reconstructed, we generate analogs by mapping the predicted BBs from SynLlama to known Enamine BBs, thereby sampling molecules from the well-defined Enamine chemical space. Under this scenario, we use nearest neighbor search algorithms with different molecular representations (SMILES and Morgan Fingerprints64) to sample Enamine neighboring BBs from the predicted BBs, as illustrated in colored inks in Figure 1(d). Since in SynLlama’s output, the RXN sequences are predicted concurrently with the BBs, our effective search space is constrained to Enamine BBs that can react through the specific RXN template. This smaller Enamine search space not only allows us to ensure the success rate of such forward syntheses but also allows us to effectively explore segments of the input molecule. Further details of the nearest neighbor search algorithms are provided in the Supplementary Information.
When constructing full synthetic pathways for reactions with multiple possible products, we select the product that most closely matches the predicted product based on SMILES string similarity. As shown in Figure 1(e), the reconstruction algorithm iteratively builds synthesis routes, utilizing all predicted BBs and RXN sequences to reconstruct or generate variations of the original molecule from the synthesizable chemical space. This reconstruction algorithm enables the SynLlama model to function as a generator for synthesizable molecules along with their corresponding synthetic pathways.